OpenStudy (anonymous):

Solve by elimination. 5x + y = 5 9x - 4y = -20

OpenStudy (saifoo.khan):

x = 0 y = 5

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

u can also use x=A^(-1)b where A^(-1) is the inverse of the coefficient matrix or row reduce [5 1 5] [9 -4 -20] 4R+R2->R2 replace row 2 with 4*row 1+ row 2 [5 1 5] [0 -29/5 -29] y = 5 5x + y = 5 => 5x = 0 x = 0